OpenStudy (anonymous):

how do you create an equation of an ellipse with a major axis of length 18 and foci located at (4,7) and (4,11)

OpenStudy (anonymous):

do you know what this looks like? we need that first

OpenStudy (anonymous):

an ellipse, but how can you tell if its horizontal or vertical?

OpenStudy (anonymous):

because of the foci the \(x\) values are the same, whereas the \(y\) values are different

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

so it is going to look like \[\frac{(y-k)^2}{a^2}+\frac{(x-h)^2}{b^2}=1\] you need the center, which is half way between \((4,7)\) and \((4,11)\) and also \(a\) and \(b\)

OpenStudy (anonymous):

you got the center?

OpenStudy (anonymous):

(4,9)

OpenStudy (anonymous):

k good so now we are here \[\frac{(y-9)^2}{a^2}+\frac{(x-4)^2}{b^2}=1\]

OpenStudy (anonymous):

how far from the center to the top part of the ellipse?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

k good, so now we know \(a=9\) and so \(a^2=81\)

OpenStudy (anonymous):

now we are here \[\frac{(y-9)^2}{9^2}+\frac{(x-4)^2}{b^2}=1\]

OpenStudy (anonymous):

how far from the center to the foci?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

b^2=4?

OpenStudy (anonymous):

got it so \(c=2\) and therefore \(c^2=4\) you need \(b^2=a^2-c^2\)

OpenStudy (anonymous):

no that was \(c\) not \(b\) we still have to find \(b\) or actually \(b^2\)

OpenStudy (anonymous):

b^2 = the square root of 77?

OpenStudy (anonymous):

i think actually \(b^2=77\) not \(b=77\)

OpenStudy (anonymous):

I get it! thank you for all the help now I understand it.

OpenStudy (anonymous):

lets check looks good http://www.wolframalpha.com/input/?i=ellipse+%28y-9%29^2%2F81%2B%28x-4%29^2%2F77%3D1 \yw